find the roots of the polynomial equation
2x^4-5x^3-17x^2+41x-21=0

Answer by stanbon(75887)

find the roots of the polynomial equation
2x^4-5x^3-17x^2+41x-21=0
-----
The coefficients add up to zero, so x = 1 is a root.
---
Use synthetic division to find other roots:
-----
1)....2....-5....-17....41....-21
......2....-3....-20....21....|..0
---------------------
Again the coefficients add up to zero, so x = 1 is again a root:
----
1)....2....-3....-20....21
......2....-1....-21...|..0
Quotient: 2x^2 -x -21
-----
Factor: 2x^2 -7x+6x -21 = 0
x(2x-7 + 3(2x-7) = 0
(2x-7)(x+3) = 0
Roots: x = 7/6, x = -3, x = 1 (multiplicity 2)
===========================
Cheers,
Stan H.
===========================